A random sample contains 23 observations. Determine the p-VALUE for each value of the calculated test statistic below and state your decision about the null hypothesis in each part by comparing the p-VALUE to ? = 0.05.

H0: ? = 50

Ha: ? ≠ 50

a) t = 1.7171

b) t=2.5083

c) t = 2.2911
Answer #1

Solution :

Given that ,

n = 23

df = 23 - 1 = 22

This is the two tailed test .

a)

P-value = 2 (t 22 > 1.7171) = 0.1

P-value > 0.05

Fail to reject the null hypothesis .

b)

P-value = 2 (t 22 > 2.5083) = 0.02

P-value < 0.05

Reject the null hypothesis .

c)

P-value = 2 (t 22 > 2.2911) = 0.0319

P-value < 0.05

Reject the null hypothesis .